OpenStudy (anonymous):

how to find value of sin 18.......plz explain in detail

OpenStudy (anonymous):

without the use of the calculator?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

how 2 derivate it

OpenStudy (anonymous):

taylor expansion can estimate it for you

OpenStudy (anonymous):

i did not get what u want to say

OpenStudy (anonymous):

sauravshakya u may try

OpenStudy (anonymous):

who is taylor

OpenStudy (anonymous):

Taylor Expansion is the way of estimating and function sinx is \[x-\frac{ x ^{3} }{ 6 }+\frac{ x ^{5} }{ 120 } -...+...\]

OpenStudy (anonymous):

5*18=90 sin(5*18)=sin90 Let x=18 then, sin(5x)=sin90

OpenStudy (anonymous):

i think u r wrog somewhere

OpenStudy (anonymous):

U can proceed from here but this will be too long but it will give u exact value.

OpenStudy (anonymous):

we dont have a formula for sin 5 thete

OpenStudy (anonymous):

From sin(5x)=1 u need to find the value of sinx

OpenStudy (anonymous):

i mean to say we dont have its expansion

OpenStudy (anonymous):

sin(5x) = sin(2x + 3x) = sin2x * cos3x + cos2x * sin3x

OpenStudy (anonymous):

hmmmmm that wd be 22 long

OpenStudy (anonymous):

5x= 90 2x+3x=90 3x=90-2x sin 3x= cos 2x now put formula of sin 3x and cos 2x in terms of sin x only.

OpenStudy (anonymous):

hmmmmmm this wd b an easy one

OpenStudy (anonymous):

18=Pi/10 sin(Pi/10)=0.3090169944 F[x]\[=x-\frac{ x ^{3} }{ 6 }+\frac{ x ^{5} }{ 120 }-... \] f[Pi/10]==0.3090170542 you decide

OpenStudy (anonymous):

Great!

OpenStudy (anonymous):

but when we will solve the quadrtic equation then we will get two values of sin theta

OpenStudy (anonymous):

which one is great

OpenStudy (anonymous):

@gohangoku58 's one

OpenStudy (anonymous):

hmmmmmmm

OpenStudy (anonymous):

because that gives exact value not approximate

OpenStudy (anonymous):

hmmmmm but two values

OpenStudy (anonymous):

which one to select

OpenStudy (anonymous):

u plz try it once

OpenStudy (anonymous):

well, sin 18 lies in first quadrant so take positive value

OpenStudy (anonymous):

hmmmmm

OpenStudy (anonymous):

good

OpenStudy (anonymous):

u'll get a cubic equation, with one root = 0 (can't be, rejected) and 2 other roots.....from which u select one with positive value..... like saura said...

OpenStudy (anonymous):

i failed 2 use common sence

OpenStudy (anonymous):

no not a cubic but a quadratic u try it

OpenStudy (anonymous):

i tried it jjust now

OpenStudy (anonymous):

sin 3x= ?

OpenStudy (anonymous):

3sinx-4 sin cube x

OpenStudy (anonymous):

3sinx-4 sin cube x<----cubic

OpenStudy (anonymous):

no but sin 2 x is also there

OpenStudy (anonymous):

*cos 2x

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

1-2sin square x

OpenStudy (anonymous):

then we can take sin in common

OpenStudy (anonymous):

no u cant 1 is there

OpenStudy (anonymous):

r u there on facebook

OpenStudy (anonymous):

who ?

OpenStudy (anonymous):

sorry for the personal Q

OpenStudy (anonymous):

we both are there...

OpenStudy (anonymous):

all of u

OpenStudy (anonymous):